# READS-V: Real-time Automated Detection of Epileptic Seizures from Surveillance Videos via Skeleton-based Spatiotemporal ViG

### v1

@@ 3.4 Partitioning spatiotemporal graph modeling
Figure 3: STViG architecture. Begin with patches extracted from the target video clip, input and positional embedding layers transform patches into feature vectors. STViG consists of four residual stages with expansion and downsampling operations. Each stage contains several proposed spatiotemporal learning layers. The output probability with dimension 1 is generated from a Sigmoid function. $N,T,C$ denotes the number of joints, frames and channels, and $H,W$ stand for the height and width of extracted patches.
Layer name Output size T STViG-Light/Base
@@
Table 1: Details of two versions of STViG. In Stem layer, $32\times 32,C$ is kernel size and output channels of Conv2D implemented on input patches ($H\times W$). Then we utilize 4-stage spatiotemporal modeling to process features. $E,C$ denotes the number of constructed edges between neighbors and output channels for spatial modeling, $1\times K_{S}\times K_{T},C$ denotes kernel size and output channels of 3D-CNN in temporal modeling. In practice, we render 3 for both $K_{S}$ and $K_{T}$.
3.5 STViG network architecture
Fig. 3 is the detailed architecture of proposed STViG network including the size of features at each layer. We keep feature size in spatial modeling, and implement downsampling at the last layer of each residual temporal block. After four residual STViG stages, we make use of global average pooling to generate a 1D feature vector, then one FC layer and sigmoid function are connected to generate output probabilities. We provide two versions of STViG models: STViG-Base and STViG-Light respectively based on different model depths and feature expansions, which are shown in Table 1.
In spatial modeling, $E=12$ and $E=2$ represent the number of edges/neighbors in inter-partition and intra-partition graph convolution, respectively. The spatial modeling phase always keeps the output channels unchanged. In temporal modeling, we set $K_{S}=3$ for simultaneously implementation of inter-partition and intra-partition temporal convolution network (TCN) operation, and set $K_{T}=3$ for considering only one frame before and after the current frame. The residual 3D-CNN block is used to expand feature dimension and downsample temporal frames.
Positional embedding. In previous studies, either skeleton-based ST-GCNs which input coordinate triplets or RGB-/Heatmap-based methods made use of specific joint coordinates or whole body context with global positional information. In this work, however, input patches in 1D sequence do not provide any positional information. Thus, we add a positional embedding to each node feature vector. Here we propose 3 different ways to inject positional information: (1) Add learnable weight $\mathbf{x}_{it}=\mathbf{x}_{it}+\mathbf{e}_{it}$; (2) Concatenation with joint coordinates $\mathbf{x}_{it}=[\mathbf{x}_{it},(x_{it},y_{it})]$; (3) Add embeddings of joint coordinates: $\mathbf{x}_{it}=\mathbf{x}_{it}+Stem(x_{it},y_{it})$;
3.6 Seizure onset decision-making
Inspired by [44], after generating output probabilities $P_{t}$ by proposed skeleton-based STViG model for video clips at time t, we can accumulate a period $\tau$ of previously detected probabilities as accumulative probabilities $AP_{t}=\sum_{i=t-\tau}^{t}P_{i}$ with detection rate $r$, then make a seizure onset decision at time $t_{onset}$ when accumulative probabilities reach a decision threshold $DT$: $t_{onset}=AP_{t}>DT$. We measure the distance between $t_{onset}$ and EEG onset as latency of EEG onset $L_{EO}$, and the distance between $t_{onset}$ and clinical onset as latency of clinical onset $L_{CO}$.
4 Experiments
4.1 Dataset
Data labeling. How to label the video clips is a challenge in video-based seizure detection. Typically, EEG-based seizure detection is treated as a binary classification task, distinguishing between interictal (pre-seizure) and ictal (during-seizure) periods. Medical experts annotate the moment from interictal to ictal periods in EEG recordings as the EEG onset. However, in video analysis, the transition from normal to seizure behavior is not as immediate. For several seconds (ranging from 1 to 30 seconds) after the EEG onset, a patient’s behavior may appear only slightly abnormal or nearly normal. We refer to this phase as the transition period. The moment when patients start exhibiting clearly abnormal actions, such as jerking, convulsion, or stiffening, is defined as clinical onset. It is straightforward to label video clips as healthy status (0) before the EEG onset and as seizure status (1) after the clinical onset. These labels (0 and 1) can be interpreted as the probability or risk of seizure onset. In terms of the transition period, it is hard to assign a precise probability to each clip. Clinically, it’s observed in the transition period that patients are more likely to appear healthy closer to the EEG onset and more likely to exhibit abnormal behaviors as the time approaches the clinical onset. To address this, we use an exponential function to assign increasing probabilities to video clips during the transition period, ranging from 0 to 1. Other options for probability-increasing functions are explored in the appendix. It is noted that the probabilities of video clips depend on the end frame of video clips lying in which period. Consequently, we define this task as a probability regression task rather than a traditional classification task.
Figure 4: Data labeling of video-based seizure detection task. For each seizure, a video recording is categorized into 3 different periods: interictal (label:0), ictal (label:1), and transition (label:0 to 1 in exponential function according to clinical phenomenon).
# pat # seizures Interictal period Transition period Ictal period
14 33 2.78h 0.16h 0.35h
Table 2: Summary of dataset used in this work.
Method Input Backbone Error FLOPs Param.
DGSTGCN [8] Skeleton GCN 16.7 0.50G 1.4M
RGBPoseConv3D [9] RGB+Skeleton 3D-CNN 13.8 12.63G 3.2M
STGCN [45] Skeleton GCN 14.1 1.22G 3.1M
MSG3D [25] Skeleton GCN 11.3 1.82G 2.7M
AAGCN [31] Skeleton GCN 9.5 1.38G 3.7M
CTRGCN [5] Skeleton GCN 8.3 0.62G 1.4M
STViG-Light (Ours) Patch ViG 6.1 0.44G 1.4M
STViG-Base (Ours) Patch ViG 5.9 1.76G 5.4M
Table 3: Comparison with previous state-of-the-arts. We reproduce state-of-the-art approaches to make comparisons. Error is averaged RMSE (%) for each video clip.
Data splitting. For each seizure event, we categorize the video segments based on their timing relative to the EEG onset and clinical onset, which are annotated by medical experts. We define the period $<2min$ after clinical onset as the ictal period, and the period $<30min$ before EEG onset as the interictal period. Given that patients’ behaviors often remain unchanged for extended periods, we strive to include a variety of actions in the extracted clips. The transition period usually ranges from $1s\sim 20s$ after EEG onset for each seizure event. In our dataset, the average transition period is approximately $18.2s$ across 33 seizures. To address the imbalance in duration between interictal and other periods, we extract clips from interictal period without overlapping, and from both transition and ictal periods with $4s$ overlappings. For fair performance comparison, we randomly extract $20\%$ clips from both transition and ictal periods from each seizure, and an equivalent number of clips (20% transition + 20% ictal) are extracted from the interictal period for validation, the rest clips are for model training. In this study, we aim to train a generalized model applicable to all patients rather than a patient-specific model. Table 2 summarize the dataset we used in this study.
Figure 5: Comparison of performance vs. efficiency.
4.2 Experimental setting
We manually set a size of $32\times 32\ (H\times W)$ for extracted patches based on the 1920$\times$ 1080 raw frame resolution, and $\sigma$ of gaussian kernel is $0.3$. All generated outputs are connected with a Sigmoid function to map the outputs as probabilities from 0 to 1, the activation function used anywhere else is chosen as ReLU, and mean square error loss function is used to train the model. During the training phase, epoch and batch size are respectively set to $200$ and $32$, we choose Adam optimizer with 1e-4 learning rate with 0.1 gamma decay every 40 epochs and 1e-6 weight decay. We save the best model with the lowest error on the validation set.
4.3 Seizure-related action recognition performance
Our STViG-Base and STViG-Light obtain $5.9\%$ and $6.1\%$ errors across all patients. We reproduce several state-of-the-art action recognition models, conducting a grid search to optimize the learning rate and weight decay for each model. We select one CNN-based approach, RGBPoseConv3d, and 5 state-of-the-art skeleton-based approaches for the comparison. According to Table 3, two STViG models both outperform previous state-of-the-arts. We also visualize model comparison of performance vs. efficiency in Fig. 5, which highlights the advantages of STViG models. Notably, STViG-Light demonstrates strong performance in both accuracy and efficiency.
The RGBPoseConv3D model is a leading RGB-based action recognition model, it was inspired by several 3D CNN architectures [38, 11, 12] and fused RGB and skeleton as input features simultaneously. Our results reveal it performs worse than skeleton-based approaches even if it contains a large number of parameters, in our opinion, its reduced effectiveness is due to the complexities of the EMU environment, which adversely affect RGB-based methods. As for skeleton-based approaches, most perform around $10\%$ error with lower FLOPs, and CTRGCN performs the best among all skeleton-based approaches. These results underscore the superiority of our skeleton-based STViG models over both RGB-based and other skeleton-based approaches in terms of accuracy and efficiency, making them highly suitable for real-time deployment.
(a) Healthy status.
(b) Head turn (seizure begins).
(c) Continue to act abnormally.
(d) Convulsion
Figure 6: Occlusion maps for model interpretability. Each image represents the end frame of the video clip. Redder partitions means more salient to the seizure-related actions.
4.4 Ablation study
We conduct ablation study with STViG-Base to evaluate the effects of positional embedding way and dynamic partition strategy on the model. In Table 4, we can see that Stem positional embedding way with dynamic partition achieves the best performance.
Effect of dynamic partition. Dynamic partition strategy shows better performance among all four types of positional embedding, which indicates that dynamic partition strategy can effectively enhance the model to learn relationships between different partitions, thereby understanding the seizure-related actions.
Effect of positional embedding. According to the results in Table 4, positional information can bring better performance for STViG model, and Stem way can outperform among three positional embedding ways. Further, we find that no positional embedding with dynamic partition can also bring a satisfactory result compared with other GCN-based approaches, which indicates that proposed STViG model has the capability to complete action recognition task without any specific or global positional information. In our point of view, the reason is that inter-partition and intra-partition representations contain enough information for action recognition no matter how these partitions are ordered or where they are.
4.5 Model interpretability
We provide interpretability with occlusion maps [48] to demonstrate the capabilities of our STViG model to learn representations of seizure-related actions between different partitions. We successively take 4 video clips of a patient from healthy status to seizure coming, as shown in Fig. 6, each image represents the end frame of corresponding clips, higher value (redder) partition means more salient to seizure-related actions, and lower value (bluer) partition means less relevant. We can see that, initially, when the patient is lying on the bed, all partitions are shown as less salient in Fig. 5(a). As the seizure begins, the patient starts to exhibit abnormal movements, such as turning the head and moving the right arm. Correspondingly, these partitions appear redder in Fig. 5(b) and Fig. 5(c), indicating their increased relevance to the seizure activity. Several seconds later, as the patient begins jerking, most partitions are shown as more salient in Fig. 5(d). This form of model interpretability provides a visual tool for doctors to efficiently analyze the progression of a seizure attack and understand the types of seizure-related actions, thereby aiding in better diagnosis and treatment planning.
Positional embedding Dynamic Partition
w/o   w.
None 16.9   9.1
Cat 14.8   8.9
Learn 13.4   8.6
Stem  8.4    5.9
Table 4: Ablation study. We utilize STViG-Base model to evaluate the effects of positional embedding way and dynamic partition. The shown values are averaged error (%) for each clip.
$DT$ FDR Latency Accumulative
w/o w.
0.5 0 $L_{EO}$ 15.3s 10.9s
$L_{CO}$2.9s -7.3s
0.2 0 $L_{EO}$ 11.4s 5.1s
$L_{CO}$6.8s -13.1s
4.6 Make seizure onset decision
After obtaining probabilities from consecutive video clips by STViG-Base model, we set up $\tau=3\ s,r=0.5\ s$ as mentioned in Section 3.6 to calculate $L_{EO},L_{CO}$ and FDR across all seizures. We also evaluate the effect of accumulative strategy on the detected latency, as shown in Table 5. We can see that performance without accumulative strategy is satisfactory based on accurate STViG model, but accumulative strategy with $DT=2$ can obtain much better performance where $5.1\ s\ L_{EO}$, -$13.1\ s\ L_{CO}$, which means seizure can be detected only $5.1\ s$ after seizure begins inside brain, and can alarm the seizure coming $13.1\ s$ before patients tend to exhibit serious convulsions. And a FDR of 0 is achieved by all conditions.
5 Conclusion
In this work, we propose a STViG model based on skeleton-based patch embeddings to recognize the epileptic patients’ actions effectively, experiments indicate proposed STViG model outperforms state-of-the-art action recognition approaches. It provides an accurate, efficient, and timely solution for the READS-V system to fundamentally help epileptic patients. In the future, we will collect more patient video data with various types of seizures to train a more generalized model for wider epileptic patients.
Supplementary Material
A. Visualization
First, we visualize the pose estimation performance comparison between our custom model and the public pre-trained model, as shown in Fig. 7, thereby demonstrating the limitation of public pre-trained model for tracking patient skeletons.
Second, we present the visualization of predictive probabilities of video clips obtained by STViG model from two different seizures, which is shown as . It is important to note that these frames are not successive, we are intended to show the clips with different predictive probabilities in a seizure. We can see that our model can detect subtle motion changes during the seizure beginning.
These two visualization results built in .gif format are included in the supplementary package.
B. Chosen of increasing probability function
In this study, we label the transition period from 0 to 1 in exponential function due to clinical phenomenon. We also choose two alternatives - linear and sigmoid functions to make a comparison with exponential function in STViG-Base model. We can see from Table 6 that the exponential labeling function outperforms the other two functions.
Linear Sigmoid Exponential
Error 6.7% 9.4% 5.2%

### v2

@@ 3.4 Partitioning spatiotemporal graph modeling
Figure 4: VSViG architecture. VSViG consists of four residual spatiotemporal (ST) ViG stages and each stage contains several proposed spatial and temporal modeling layers. $N,T,C$ denotes the number of joints, frames and channels, and $H,W$ stand for the height and width of extracted patches.
3.5 VSViG network architecture
Fig. 4 shows the detailed structure of proposed VSViG network including the shape of features at each layer. We keep feature size in spatial modeling, and implement downsampling at the last layer of each residual temporal block. After four residual spatiotemporal ViG stages, we make use of global average pooling to generate a 1D feature vector, then one FC layer and sigmoid function are connected to generate output probabilities. In spatial modeling, $E=12$ and $E=2$ represent the number of edges (neighbors) in inter-partition and intra-partition graph convolution, respectively. The spatial modeling phase always keeps the output channels unchanged. In temporal modeling, we set $K_{S}=3$ for simultaneous implementation of inter-partition and intra-partition temporal convolution network (TCN) operation, and set $K_{T}=3$ for considering only one frame before and after the current frame. The residual 3D-CNN block is used to expand feature dimension and downsample temporal frames.
We propose two versions of VSViG models: VSViG and VSViG-Light respectively based on different model depths and feature expansions, they are designed for accurate or efficient purposes. Their details are illustrated in Appendix.
Positional embedding. In previous studies, either skeleton-based STGCNs which input coordinate triplets or RGB-/Heatmap-based methods made use of specific joint coordinates or whole body context with global positional information. In this work, however, input patches in 1D sequence do not provide any positional information. Thus, we add a positional embedding to each node feature vector. Here we propose 3 different ways to inject positional information: (1) Add learnable weight: $\mathbf{x}_{it}=\mathbf{x}_{it}+\mathbf{e}_{it}$; (2) Concatenation with joint coordinates: $\mathbf{x}_{it}=[\mathbf{x}_{it},(x_{it},y_{it})]$; (3) Add embeddings of joint coordinates: $\mathbf{x}_{it}=\mathbf{x}_{it}+Stem(x_{it},y_{it})$;
3.6 Seizure onset decision-making
Inspired by [49], after generating output probabilities $P_{t}$ by proposed skeleton-based VSViG model for video clips at time t, we can accumulate a period $\tau$ of previously detected probabilities as accumulative probabilities $AP_{t}=\sum_{i=t-\tau}^{t}P_{i}$ with detection rate $r$, then make a seizure onset decision at time $t_{onset}$ when accumulative probabilities reach a decision threshold $DT$: $t_{onset}=AP_{t}>DT$. We measure the distance between $t_{onset}$ and EEG onset as latency of EEG onset $L_{EO}$, and the distance between $t_{onset}$ and clinical onset as latency of clinical onset $L_{CO}$.
Figure 5: Data labeling for a regression-based task. For each seizure, a video recording is categorized into 3 different periods: interictal (label: 0), ictal (label: 1), and transition (label: 0 to 1 in exponential function according to clinical phenomenon).
4 Experiments
4.1 Dataset
We acquire surveillance video data of epileptic patients from the local hospital, this dataset includes 14 epileptic patients with 33 focal or generalized tonic-clonic seizures. As for each seizure, we utilize all ictal periods, and 5-30 min interictal periods before the EEG onset. We segment the raw video recording into video clips with 5 s for model training and manually label the video clips as likelihood for regression task according to Fig. 5. Naive binary classification of healthy and unhealthy status cannot accurately recognize the early occurrence of seizures, the introduced increasing probability way is consistent with the likelihood of real seizure attack risk. And an exponential way is more similar to the clinical phenomenon that the closer time to the clinical onset, the higher risk is obtained. We also test other functions to build increasing probability labels for this regression task, which is shown in Appendix.
Eventually, 2.78 h interictal period, 0.16 h transition period and 0.35 h ictal period video are obtained for the model training, and overlapping extraction operation is utilized to overcome the issue of data imbalance. Then we split 70% of each period for training, and the rest 30% for validation. More details information about the dataset is illustrated in Appendix.
Performance metric. In this paper, we define a regression task intended for early seizure detection, RMSE metric measured between true video clip and predictive video clip is used to compare the model performance.
Data usage. The data collection from hospital is carefully conducted and the use of data was approved by hospital Institutional Review Board. We will release this dataset when all sensitive and private information is removed. Code and fine-tuned pose estimation model will be also released.
4.2 Experimental setting
We manually set a size of $32\times 32\ (H\times W)$ for extracted patches based on the 1920$\times$ 1080 raw frame resolution, and $\sigma$ of gaussian kernel is $0.3$. All generated outputs are connected with a Sigmoid function to map the outputs as probabilities from 0 to 1, the activation function used anywhere else is chosen as ReLU, and mean square error loss function is used to train the model. During the training phase, epoch and batch size are respectively set to $200$ and $32$, we choose Adam optimizer with 1e-4 learning rate with 0.1 gamma decay every 40 epochs and 1e-6 weight decay.
4.3 Seizure-related action recognition performance
According to Table 1, VSViG and VSViG-Light respectively obtain $5.9\%$ and $6.1\%$ errors across all patients, which outperform all previous state-of-the-art approaches. Furthermore, VSViG-Light only obtains 0.44G FLOPs and 1.4M model size which is desirable for real-time deployment.
We also reproduce several state-of-the-art action recognition and video-based seizure detection models to make comparisons with proposed VSViG model. There are four selective state-of-the-art video-based seizure detection approaches, they are all RGB-based strategies utilizing a CNN-based model to extract features that are followed by a sequence-based model. In terms of skeleton-based action recognition models, several state-of-the-art GCN-based models are chosen, and one RGB+skeleton fusion method is also considered.
Table 1: Model comparisons. We reproduce state-of-the-art approaches to make comparisons with proposed VSViG. (OF: optical flow, $\clubsuit$: Video-based seizure detection models, $\spadesuit$: Skeleton-based action recognition models, $\bigstar$: This work)
Method Input RMSE(%) GFLOPs Param.(M)
$\clubsuit$ I3D+LSTM[22] RGB 18.7 57.3 28.0
$\clubsuit$ CNN+LSTM[51] RGB 17.7 5.0 17.1
$\clubsuit$ R3D+LSTM[34] RGB 16.5 163.1 31.9
$\clubsuit$ R3D+ViT[30] RGB 14.2 163.1 32.6
$\spadesuit$ DGSTGCN [10] Skeleton 16.7 0.50 1.4
$\spadesuit$ STGCN [50] Skeleton 14.1 1.22 3.1
$\spadesuit$ RGBPoseConv3D [11] RGB+Skeleton 13.8 12.63 3.2
$\spadesuit$ JOLOGCN [3] OF in Patches 12.3 37.30 6.9
$\spadesuit$ MSG3D [29] Skeleton 11.3 1.82 2.7
$\spadesuit$ AAGCN [36] Skeleton 9.5 1.38 3.7
$\spadesuit$ CTRGCN [7] Skeleton 8.3 0.62 1.4
$\spadesuit$ HDGCN [24] Skeleton 7.9 1.60 1.7
$\bigstar$ VSViG-Light Patches 6.1 0.44 1.4
$\bigstar$ VSViG Patches 5.9 1.76 5.4M
We can see that RGB-based approaches perform worse in both error and FLOPs than skeleton-based approaches, it is consistent with RGB-based approach’s drawbacks: large model size and can easily be disturbed by nuisances. Furthermore, several previous video-based seizure detection model rely on the pre-trained RGB-based 3D CNN model to extract features, which extremely increases the computational resources. As for skeleton-based approaches, most perform around $10\%$ error with lower FLOPs, and CTRGCN performs the best among all skeleton-based approaches. These results underscore the superiority of our skeleton-based patch embedding VSViG models over both RGB-based and skeleton-based GCN approaches
4.4 Visualization
We visualize the pose estimation performance comparison between our fine-tuned model and the public pre-trained model, as shown in Fig. 6, thereby demonstrating the limitation of public pre-trained model for tracking patient skeletons. Secondly, we present the visualization of predictive probability of each video clip obtained by VSViG model in a successive seizure from two different seizures, which is shown as Fig. 7.
Figure 6: Pose estimation performance comparison between public pre-trained model and fine-tuned model. Black circles are used to mask sensitive information.
Figure 7: Predictive probabilities/likelihoods of video clips from two different seizures are obtained by VSViG model. Black boxes are used to keep privacy.
4.5 Ablation study
We conduct ablation studies with VSViG to evaluate the effects of several modules and settings on the model performance, all shown values in Table 2 are RMSE errors on the validation set.
Effect of spatial modeling module. There are several spatial modeling modules in proposed VSViG, we conduct an ablation study here to prove the effectiveness of inter-partition and intra-partition modules. We can see from Table 2(a) that proposed inter-intra module outperforms other conditions. Inter-partition and intra-partition fundamentally extract coarse and fine features respectively, and Inter-Intra strategy is consistent with coarse-to-fine feature extraction that can learn the representations better.
Effect of temporal modeling kernel size. As for the temporal modeling, a kernel size of $(3,3)$ stands for $K_{S}$, $K_{T}$ in TCN operation, $K_{S}=3$ means spatial neighbor joints are also considered, meanwhile $K_{S}=1$ is not. And we also compare 3 and 5 convolution size in temporal dimension $K_{T}$. The result in Table 2(b) shows that considering neighbor nodes and $K_{T}=3$ performs best.
Effect of dynamic partition. In Table 2(c), we can see that Stem positional embedding way with dynamic partition achieves the best performance. Dynamic partition strategy shows better performance, which indicates it can effectively enhance the model to learn relationships between different partitions, thereby understanding the seizure-related actions.
Effect of positional embedding. Also, positional information can bring better performance for VSViG model, and Stem way outperforms all conditions. Further, no positional embedding with dynamic partition can also achieve a satisfactory result compared with other GCN-based approaches that utilize coordinates as input, this indicates that proposed VSViG model has the capability to complete an action recognition task without any specific or global positional information. In our point of view, the reason is that inter-partition and intra-partition representations contain enough information for action recognition no matter how these partitions are ordered or where they are.
Effect of gaussian kernel size. During the phase of fine-tuning lightweight-openpose, $\sigma$ size should be set in advance. Smaller size brings less RGB information in extracted patches, meanwhile larger size brings more contextual nuisances. In this ablation study Table 2(d), 0.3 relative to the patch size is the best. It is noted that the chosen patch size depends on the video resolution and size of patient in the video, $32\times 32$ is suitable for this scenario.
Table 2: Ablation studies on VSViG. All experiments are conducted with VSViG version, and values are RMSE.
(a) Spatial modeling module.
 Inter  Intra  RMSE
 ✗  ✗  13.0
 ✓  ✗  11.7
 ✗  ✓  9.2
 ✓  ✓ 5.9
(b) Temporal modeling kernel size.
$(K_{S},K_{T})$  RMSE
 (1, 3)  9.9
 (1, 5)  10.8
 (3, 3) 5.9
 (3, 5)  6.5
(c) Positional embedding and dynamic partition.
Positional embedding Dynamic partition
w/o   w.
None 16.9   9.1
Cat. 14.8   8.9
Learnable 13.4   8.6
Stem (x, y)  8.4    5.9
(d) Gaussian kernel ($\sigma$) size.
$\sigma$  RMSE
 0.1  10.1
 0.2  8.5
 0.3 5.9
 0.4  8.6
 0.5  7.6
Figure 8: Occlusion maps for model interpretability. Each image represents the end frame of the video clip. Redder partitions mean more salient to the seizure-related actions. (a)-(d) present a patient suffering a seizure from healthy status to convulsion.
4.6 Model interpretability
We provide interpretability with occlusion maps [53] to demonstrate the capabilities of our VSViG model to learn representations of seizure-related actions within different partitions. We successively take 4 video clips of a patient from healthy status to convulsion, as shown in Fig. 8, each image represents the end frame of corresponding clips, higher value (redder) means the partition is more salient to seizure-related action, and lower value (bluer) means less relevant. We can see that, initially, when the patient is lying on the bed, all partitions are shown as less salient in Fig. 8(a). As the seizure begins, the patient starts to exhibit abnormal movements, such as turning the head and moving the right arm. Correspondingly, these partitions appear redder in Fig. 8(b) and Fig. 8(c), indicating their increased relevance to the seizure activity. Several seconds later, as the patient begins jerking, most partitions are shown as more salient in Fig. 8(d).
4.7 Make seizure onset decision
After obtaining probabilities from consecutive video clips by VSViG model, we set up $\tau=3\ s,r=0.5\ s$ as mentioned in Section 3.6 to calculate $L_{EO},L_{CO}$ and FDR across all seizures. We also evaluate the effect of accumulative strategy on the detected latency, as shown in Table 3. We can see that performance without accumulative strategy is satisfactory based on the accurate VSViG model, but accumulative strategy with $DT=0.3$ can obtain much better performance where $5.1\ s\ L_{EO}$, -$13.1\ s\ L_{CO}$, which means seizure can be detected only $5.1\ s$ after seizure begins inside brain, and can alarm the seizure coming $13.1\ s$ before patients tend to exhibit serious convulsions. And FDR of 0 is also obtained. The specific results of every patient are shown in Appendix.
$DT$ FDR Latency Accumulative
w/o w.
0.3 0 $L_{EO}$ 11.4s 5.1s
$L_{CO}$6.8s -13.1s
5 Conclusion
In this work, we proposed a VSViG model utilizing skeleton-based patch embeddings to recognize the epileptic patients’ actions effectively, experiments indicate proposed VSViG model outperforms state-of-the-art action recognition approaches. VSViG provides an accurate, efficient, and timely solution for video-based early seizure detection to fundamentally help epileptic patients.
In the future, we will collect more patient video data with various types of seizures to train a more generalized model for a large number of epileptic patients. And action recognition-based VSViG model can also be extended to detect other movement-based diseases, e.g. Parkinson’s disease, fall detection.
Acknowledgement
This work was supported by STI2030-Major Projects (2022ZD0208805), ”Pioneer” and ”Leading Goose” R&D Program of Zhejiang (2024C03002), and the Key Project of Westlake Institute for Optoelectronics (Grant No. 2023GD004).
A. Data details
Data labeling. How to label the video clips is a challenge in video-based early seizure detection. Typically, EEG-based seizure detection is treated as a binary classification task, distinguishing between interictal (pre-seizure) and ictal (during-seizure) periods. Medical experts annotate the moment from interictal to ictal periods in EEG recordings as the EEG onset. However, in video analysis, the transition from normal to seizure behavior is not as immediate. For several seconds (ranging from 1 to 30 seconds) after the EEG onset, a patient’s behavior may appear only slightly abnormal or nearly normal. We refer to this phase as the transition period. The moment when patients start exhibiting clearly abnormal actions, such as jerking, convulsion, or stiffening, is defined as clinical onset. It is straightforward to label video clips as healthy status (0) before the EEG onset and as seizure status (1) after the clinical onset. These labels (either 0 or 1) can be interpreted as the probability or risk of seizure onset. In terms of the transition period, it is hard to assign a precise probability to each clip. Clinically, it’s observed in the transition period that patients are more likely to appear healthy closer to the EEG onset and more likely to exhibit abnormal behaviors as the time approaches the clinical onset. To address this, we use an exponential function to assign increasing probabilities to video clips during the transition period, ranging from 0 to 1. Other options, sigmoid and linear functions, are compared to exponential way in Table 4. It is noted that the probabilities of video clips depend on the end frame of video clips lying in which period. Consequently, we define this task as a probability regression task rather than a traditional classification task.
Linear Sigmoid Exponential
RMSE 6.7% 9.4% 5.9%
Data splitting. For each seizure event, we categorize the video segments based on their timing relative to the EEG onset and clinical onset, which are annotated by medical experts. We define the period $<2min$ after clinical onset as the ictal period, and the period $<30min$ before EEG onset as the interictal period. Given that patients’ behaviors often remain unchanged for extended periods, we strive to include a variety of actions in the extracted clips. The transition period usually ranges from $1s\sim 20s$ after EEG onset for each seizure event. In our dataset, the average transition period is approximately $18.2s$ across 33 seizures. To address the imbalance in duration between interictal and other periods, we extract clips from interictal period without overlapping, and from both transition and ictal periods with $4s$ overlappings. For fair performance comparison, we randomly extract $20\%$ clips from both transition and ictal periods from each seizure, and an equivalent number of clips (20% transition + 20% ictal) are extracted from the interictal period for validation, the rest clips are for model training. In this study, we aim to train a generalized model applicable to all patients rather than a patient-specific model.
B. Detailed architecture of VSViG model
In Table 5, we present detailed architecture of proposed VSViG model and its lighten version.
Table 5: Details of two versions of VSViG. In Stem layer, $32\times 32,C$ is kernel size and output channels of Conv2D implemented on input patches ($H\times W$). Then we utilize 4-stage spatiotemporal modeling to process features. $E,C$ denotes the number of constructed edges between neighbors and output channels for spatial modeling, $1\times K_{S}\times K_{T},C$ denotes kernel size and output channels of 3D-CNN in temporal modeling.
Layer name Output size T VSViG-Light/VSViG